# Improving Spiking Neural Network Accuracy With Color Model Information Encoded Bit Planes

### v1

@@ VI-D Experiment setup
with $l_{n}$ is loss of each mini-batch, $x$ is the input, $y$ is the target, $w$ is the weight, total amount of class $C$ where mini-class $c,i\in C$.
VII Results
@@
In this section we will discuss results of our main experiments along with some interesting findings beside original hypotheses.
VII-A The effect of bit plane encoding on colored images
@@
For our experiment, we conducted tests on various image classification datasets, including CIFAR10[41], CIFAR100[41], EuroSAT[46], Caltech101[47], Caltech256[48], Imagenette[49] (which is a subset of ImageNet[50]), GTSRB[51], and Food101[52]. We allocated 80% of each dataset for training and the remaining 20% for validation. All datasets were processed with a batch size of 16 images for 100 epochs on an NVIDIA RTX 3090 GPU, where CIFAR10[41], CIFAR100[41] and EuroSAT[46] are trained on original input size while input of all other dataset are resize to 224x224 pixels.
As depicted in Table II, the average accuracy and loss curves across all tested color models demonstrated improved performance. Despite a deterministic increase in inference time and input size, the results indicate that the information extracted from color model bit planes enhanced the training process. This enhancement serves as an information supplement, leading to higher performance overall without increasing the model size.
Accuracy of encoding methods (%)
Dataset Poisson rate coding With RGB encoded (Ours) With CMY encoded (Ours) With YCbCr encoded (Ours) With HSL encoded (Ours) With HSV encoded (Ours) With CIE XYZ encoded (Ours) With CIE LAB encoded (Ours)
CIFAR10[41] 70.69 73.49 73.06 73.64 73.39 73.06 73.63 74.06
CIFAR100[41] 38.57 42.15 41.31 41.62 41.94 41.91 41.73 42.03
Caltech101[47] 61.67 64.55 65.65 65.53 66.46 68.41 65.30 66.22
Caltech256[48] 30.04 41.51 40.65 40.84 40.55 39.94 40.50 40.79
EuroSAT[46] 84.39 88.48 86.61 88.11 88.74 88.81 88.87 87.48
Imagenette[49] 72.99 78.68 77.35 78.78 78.55 78.50 77.35 78.70
Food101[52] 19.62 36.95 37.71 36.98 38.03 36.41 37.10 35.87
GTSRB[51] 93.36 95.35 51.36 70.68 79.18 80.86 51.78 51.78
TABLE II: Comparison of best validation accuracy in percentage between traditional rate coding method and our encoding methods, best performing model is emphasized.
VII-B The effect of bit plane encoding on low resolution grayscale images
Accuracy of encoding methods (%)
Dataset Poisson rate coding with time step T=10 Poisson rate coding with time step T=10 and combined RGB bit plane encoding RGB bit plane encoding with no rate coding
MNIST[53] 98.43 98.51 97.63
Fashion-MNIST[54] 87.99 88.78 85.76
KMNIST[55] 93.85 94.60 91.39
TABLE III: Comparison of best validation accuracy in percentage between encoding methods, best performing model is emphasized.
Our encoding scheme is designed to leverage color model information to enhance the performance of SNN models, which is why testing on grayscale images was not initially prioritized. However, given that bit-plane encoding can be feasibly applied to single-channel grayscale images, we aim to investigate whether this approach yields any significant or insightful results. Although grayscale images consist of a pixel intensity range between [0, 255], we can adapt our encoding process by applying RGB-based encoding with a reduced number of channels. This method would produce eight bit-plane channels that represent the image information in binary form.
As demonstrated in Table III (for further details see Figure 5 in Appendix References), the combination of Poisson rate coding with bit-plane encoding outperformed other coding methods in terms of validation accuracy. Notably, using only bit-plane information resulted in only a slight reduction in accuracy. As discussed in the previous section, we hypothesize that although bit-plane encoding contains less information compared to traditional rate coding, the model can still converge to a certain extent in computer vision tasks. We further suspect that this performance deficit may be less pronounced in less complex tasks (in this situation, is a low resolution grayscale image recognition task).
VIII Conclusion
@@
In conclusion, our research introduces a novel encoding method for SNN that leverages bit planes derived from various color models of input image data. This approach prove to enhances the computational accuracy of SNN without increasing the model size. Through extensive experimental validation, we have demonstrated the effectiveness of our encoding strategy across multiple computer vision tasks, achieving notable performance gains compared to conventional methods. This study is pioneering in its application of color spaces within the context of spiking neural networks, and we believe that our findings open new avenues for the development of more efficient and effective SNN models. Our work sets a foundation for future research and applications, potentially advancing the field of computational neuroscience and artificial intelligence.

### v2

@@ V-A Experimental settings
where $l_{n}$ is loss of each mini-batch, $x$ is the input, $y$ is the target, $w$ is the weight, $C$ is the total amount of class where $i\in C$ and $c$ is the mini-classes.
In the first part of our experiments, we only use grayscale image datasets to clearly see the effects of bit plane coding. In the second part, popular datasets of color images are used to see the practical effects of both bit plane coding and color models. We allocated 80% of each dataset for training and the remaining 20% with deterministic random seeding for validation. All datasets were processed with a batch size of 16 images for 100 epochs on an NVIDIA RTX 3090 GPU.
V-B Experimental Results
@@
In this part, we will discuss results of our main experiments along with some interesting findings beside original hypotheses.
V-B1 The effect of bit plane coding on grayscale images
Dataset Accuracy of coding methods (%)
Baseline Bit plane coding only Proposed method
MNIST[36] 98.62 97.87 98.93
KMNIST[46] 94.13 92.33 95.58
Fashion-MNIST[47] 88.83 86.66 90.19
TABLE III: Comparison of validation accuracy among three coding methods. Best value is highlighted in bold.
Dataset Analysis results of coding methods (bits)
Baseline Bit plane coding only Proposed method
AKC SE AKC SE AKC SE
MNIST[36] 721.5 0.53 772.8 0.56 744.4 0.55
KMNIST[46] 1093 0.70 1221 0.70 1150 0.70
Fashion-MNIST[47] 1384 0.86 1377 0.82 1381 0.84
TABLE IV: Comparison of analysis results among coding methods, contains the approximated Kolmogorov complexity (abbreviated AKC) and Shannon Entropy (abbreviated SE). Highest values of AKC are denoted in red and lowest are denoted in blue.
First, we aim to investigate whether the proposed approach yields any significant or insightful results without impacts of color information. For that, we employed three popular datasets of grayscale images, including MNIST[36], KMNIST[46], Fashion-MNIST[47]. The results of MNIST dataset were already shown in Table I. However, they are still included here to see the behaviors of the coding methods over different datasets. As grayscale images consist of only intensity values ranging between [0, 255], we apply our coding process with one channel. This process would produce eight bit-plane channels that represent the image information in binary form. In this part, we investigate three options, namely 1) the baseline model (spike coding only), 2) using bit plane coding only, 3) the proposed model (spike coding combined with bit plane coding).
As demonstrated in Table III, the combination of spike coding with bit-plane coding outperformed other coding methods in terms of validation accuracy. As mentioned, using only bit-plane coding resulted in a reduction in accuracy; however, the combination of spike coding with bit plane coding (i.e. the proposed method) provides gain in all datasets. For any method, the accuracy value is highest with the MNIST dataset and lowest with the Fashion-MNIST dataset. This could be because among these three datasets, MNIST is the simplest one while Fashion-MNIST is the most complex one. Interestingly, the gain of the proposed method compared to the baseline is increased from 0.31% (with MNIST) to roughly 1.4% (with KMNIST and Fashion-MNIST). That means the impact of bit plane coding is more prominent with more complex datasets.
Upon further analysis of spike distribution complexities for various coding methods, as presented in Table IV, several insights emerge.
First, we observe that the baseline method has the lowest complexity in MNIST and KMNIST datasets, while it has the highest complexity in the Fashion-MNIST dataset. We suspect that this difference stems from the nature of these datasets. Both MNIST and KMNIST primarily use binary pixel values, with zeros representing the background and ones indicating the presence of a character. This binary representation results in a simpler spike distribution. In contrast, Fashion-MNIST employs a wider range of pixel values to capture the textures of various clothing items, which introduces greater complexity into the spike distribution when using baseline coding methods. This observation suggests that more complex visual datasets may lead to even more pronounced differences in spike distribution when using these coding methods.
Additionally, by introducing a hybrid coding approach that combines baseline and bit plane coding methods, we hypothesize the potential existence of an optimal complexity threshold at which the SNN might achieve peak performance. This idea of an optimizable global maximum suggests that complexity thresholds could play a key role in determining SNN effectiveness. Based on our data, we propose that measuring spike distribution complexity could be crucial in tuning SNNs to perform optimally across different datasets and tasks.
V-B2 The effect of bit plane coding on color images
Dataset Accuracy of coding methods (%)
Baseline Bit plane coding only Proposed method
CIFAR10[26] 70.69 37.28 73.49
CIFAR100[26] 38.57 11.28 42.15
Caltech101[48] 61.67 43.28 64.55
Caltech256[49] 30.04 21.75 41.51
EuroSAT[50] 84.39 79.87 88.48
Imagenette[51] 72.99 48.20 78.68
Food101[52] 19.62 6.29 36.95
GTSRB[53] 93.36 95.85 95.35
TABLE V: Average validation accuracy on various computer vision dataset of our proposed method.
Dataset Analysis results of coding methods (bits)
Baseline Bit plane coding only Proposed method
AKC SE AKC SE AKC SE
CIFAR10[26] 7530 0.995 6764 0.996 7190 0.995
CIFAR100[26] 7374 0.995 6710 0.996 7079 0.995
Caltech101[48] 310k 0.993 260k 0.992 279k 0.992
Caltech256[49] 307k 0.991 257k 0.990 276k 0.990
EuroSAT[50] 30.3k 0.953 22.9k 0.944 27.0k 0.949
Imagenette[51] 334k 0.987 282k 0.994 302k 0.992
Food101[52] 339k 0.989 288k 0.995 308k 0.993
GTSRB[53] 308k 0.902 2914 0.223 5791 0.019
TABLE VI: Comparison of analysis results among coding methods, contains the approximated Kolmogorov complexity (abbreviated AKC) and Shannon Entropy (abbreviated SE), values in thousands are represented with k (e.g., 22k = 22,000). Highest values of AKC are denoted in red and lowest are denoted in blue.
Dataset Analysis results of coding methods (bits)
With RGB encoded (Ours) With CMY encoded (Ours) With YCbCr encoded (Ours) With HSL encoded (Ours) With HSV encoded (Ours) With CIE XYZ encoded (Ours) With CIE LAB encoded (Ours)
AKC SE AKC SE AKC SE AKC SE AKC SE AKC SE AKC SE
CIFAR10[26] 7190 0.995 7086 0.992 6671 0.972 6634 0.927 6752 0.957 7083 0.986 6857 0.988
CIFAR100[26] 7079 0.995 6975 0.991 6606 0.974 6551 0.929 6660 0.955 6984 0.987 6769 0.985
Caltech101[48] 279k 0.992 269k 0.978 231k 0.957 241k 0.905 242k 0.929 268k 0.989 239k 0.971
Caltech256[49] 276k 0.990 265k 0.976 233k 0.957 243k 0.907 244k 0.933 244k 0.989 242k 0.972
EuroSAT[50] 27.0k 0.949 26.5k 0.968 25.2k 0.944 25.3k 0.993 25.8k 0.950 26.2k 0.932 25.9k 0.966
Imagenette[51] 302k 0.992 292k 0.989 267k 0.957 271k 0.901 276k 0.941 288k 0.979 277k 0.976
Food101[52] 308k 0.993 297k 0.990 273k 0.986 277k 0.908 280k 0.927 294k 0.975 281k 0.969
GTSRB[53] 5791 0.019 5706 0.006 17.6k 0.489 86.1k 0.462 92.9k 0.499 5704 0.005 5703 0.006
TABLE VII: Comparison of analysis results with our proposed coding method across all tested color models, contains the approximated Kolmogorov complexity (abbreviated AKC) and Shannon Entropy (abbreviated SE) measured in bits, values in thousands are represented with k (e.g., 22k = 22,000). Highest values of AKC are denoted in red and lowest are denoted in blue.
Dataset Accuracy of coding methods (%)
With RGB encoded (Ours) With CMY encoded (Ours) With YCbCr encoded (Ours) With HSL encoded (Ours) With HSV encoded (Ours) With CIE XYZ encoded (Ours) With CIE LAB encoded (Ours)
CIFAR10[26] 73.49 73.73 73.64 73.39 73.06 73.63 74.06
CIFAR100[26] 42.15 41.31 41.62 41.94 41.91 41.73 42.03
Caltech101[48] 64.55 65.65 65.53 66.46 68.41 65.30 66.22
Caltech256[49] 41.51 40.65 40.84 40.55 39.94 40.50 40.79
EuroSAT[50] 88.48 86.61 88.11 88.74 88.81 88.87 87.48
Imagenette[51] 78.68 77.35 78.78 78.55 78.50 77.35 78.70
Food101[52] 36.95 37.71 36.98 38.03 36.41 37.10 35.87
GTSRB[53] 95.35 51.36 70.68 79.18 80.86 51.78 51.78
Average 65.15 59.21 62.02 63.36 63.49 59.53 59.61
TABLE VIII: Comparison of validation accuracy among coding methods.
@@ V-B2 The effect of bit plane coding on color images
For our experiment, we conducted tests on various color-image classification datasets, including CIFAR10[26], CIFAR100[26], EuroSAT[50], Caltech101[48], Caltech256[49], Imagenette[51] (which is a subset of ImageNet[54]), GTSRB[53], and Food101[52]. The default color model of these datasets is RGB. Due to computing resource constraints of the simulation tool, large images are downscaled to 224x224 pixels. So, among these datasets, CIFAR10[26], CIFAR100[26] and EuroSAT[50] are trained with original input size while images of all other dataset are resized to 224x224 pixels.
Across all tested datasets, our proposed methods consistently outperform the baseline method, with the notable exception of the GTSRB[53] dataset, where bit-plane coding alone appears to yield the best performance without the need for combination with traditional spike coding. This observation suggests that while our methods generally enhance performance by improving the representation of spike distributions, certain datasets may favor different coding approaches depending on their visual complexity and specific characteristics. In the GTSRB[53] case, it seems that bit-plane coding alone is sufficient to capture the essential features for effective processing.
The analysis results presented in Table VI highlight the relationship between complexity and performance in different coding methods. Bit-plane coding consistently exhibits lower complexities across datasets, while traditional spike coding tends to produce higher complexity values. This outcome aligns with our hypothesis that complexity increases in visual datasets with greater detail and variation.
In further investigating the GTSRB[53] dataset, we observed a significant reduction in spike distribution complexity (over 99% when comparing baseline with proposed scheme and pure bit planes coding in Table VI) alongside improved model performance. Usually a decrease in complexity can sometimes indicate a loss of critical information, it may also suggest that the spike distribution has become more streamlined, capturing essential features more effectively. However, a complexity reduction at such large scale could also introduce elements of instability (shown in Table VII), as it can be challenging to discern whether the decreased complexity represents a beneficial simplification or a detrimental loss of information.
Consequently, while our findings indicate promising results, we suggest further exploration of complexity thresholds to ensure consistent performance across varying datasets.
V-B3 Impacts of color models
In this part, also we investigate two options, namely 1) the baseline method and 2) proposed method.
The accuracy results of different cases are shown in Table VIII. We can see that, when using RGB color model, the accuracy of the proposed method is always higher than that of the baseline. The gain ranges from about 2% (with GTSRB[53] dataset) up to 11% (with Caltech256[49] dataset), depending on the employed dataset. This suggests that the benefit of the proposed method when applying to color images is more significant than when applying to grayscale images.
Interestingly, the results show that RGB is not always the best color model. Among the eight datasets, RGB provides the best performance in three cases (namely CIFAR100[26], Caltech256[49], GTSRB[53]). CMY does not achieve any best results, whereas each of the other five color models achieves the best performance once. Especially, with Caltech101[48] dataset, the accuracy of HSV color model is 68.41%, which is much higher than that of RGB color model (64.55%).
When considering the average of accuracy values across all datasets, the RGB color model is significantly better than other color models (mentioned the last row of Table VIII) and even though there are performance degradation on GTSRB[53], averaged result across all dataset proven to at least out performed baseline method.
From the above discussion, the key findings of our research can be summarized as follows.
  • The proposed method provides benefits for both grayscale images and color images.
  • The gain of the proposed method depends on the employed dataset. For practical color datasets, the gain is from 2% up to 11%.
  • In terms of average accuracy value over all datasets, RGB is the best color model and CMY is the worst one.
  • However, considering individual datasets, RGB is not always the best color model. It provides the best performance in only three datasets.
  • Especially, with the Caltech101 dataset, the performance of HSV color model (68.41%) is much better than that of the RGB color model (64.55%).
VI Discussion
VI-A Complexity or entropy?
In this study, we employ two primary metrics, an approximated Kolmogorov complexity and Shannon entropy, to analyze the effects of bit-plane coding on the distribution of spike trains. These metrics were chosen because they are widely used in signal processing and information theory, making them well-suited for analyzing the coded spike and bit-plane signals. By treating both coded spikes and bit planes as types of signals, we believe that Kolmogorov complexity and Shannon entropy can provide valuable insights into the complexity of spike distributions. While we conducted limited analysis with Shannon entropy, our results indicate a positive correlation between Shannon entropy and Kolmogorov complexity. This suggests that Shannon entropy could serve as an alternative to Kolmogorov complexity for those interested in measurement and analysis. However, we find Kolmogorov complexity to be more effective, as it provides clearer distinctions between measurements when compared to Shannon entropy.
Despite achieving substantial insights through these metrics, we recognize that measuring complexity and entropy alone may not fully capture the intricacies of image data. Images inherently contain structural complexity and spatial information, which can be difficult to quantify and analyze. Focusing solely on spike signals as sequential data may overlook these aspects. A helpful analogy is the historical debate in physics regarding the nature of light, with some scientists viewing light as a wave and others as particles. It was only with the discovery of wave-particle duality through experiments like the double-slit experiment that a more comprehensive understanding emerged. Similarly, viewing spike signals solely as sequential signals could reveal certain patterns but may not capture the full complexity underlying the optimization process in image processing tasks. Currently, tools designed specifically for spike signal analysis that incorporate spatial information for image-based tasks are underexplored. This gap highlights the need for further research into analysis methods that can integrate both signal and spatial information for more complete insights.
VII Conclusion
@@
In conclusion, our research introduces a novel coding method for SNN that leverages bit planes derived from various color models of input image data. This approach prove to enhances the computational accuracy of SNN without increasing the model size. Through extensive experimental validation, we have demonstrated the effectiveness of our coding strategy across multiple computer vision tasks, achieving notable performance gains compared to conventional methods. This study is pioneering in its application of color models within the context of spiking neural networks, and we believe that our findings open new avenues for the development of more efficient and effective SNN models. Our work sets a foundation for future research and applications, potentially advancing the field of computational neuroscience and artificial intelligence.
Appendix A CMY or CMYK?
In the initial stages of our experimental setup and algorithm design, we encountered challenges in selecting the appropriate device-dependent color model for our experiments. Specifically, we debated whether to use the CMY or CMYK color model. Each has advantages in different applications, but their compatibility with the neural networks used in our research required careful consideration. In image processing and printing, the CMYK model was developed to improve print quality by reducing ink usage in darker regions. This is achieved by adding a K component, which reduces the CMY values proportionally, leading to enhanced contrast and more economical ink usage in printed images.
However, when adapting this approach for digital image processing, particularly for direct usage on digital devices, the inclusion of a K channel becomes less beneficial. Unlike printed images, digital displays do not require ink density adjustments, making the K channel unnecessary for accuracy. For our experiment, this would introduced dimensional conflicts in tested SNN architectures and coding schemes since they are previously optimized for three-channel data inputs (RGB), including a fourth channel would require additional modifications to model architecture, potentially affecting performance and comparability.
Furthermore, in our image processing pipeline, we observed that converting images to the CMYK color model using standard image processing libraries (e.g., PIL[55]) did not apply undercolor removal during conversion to CMYK model, resulting in a K channel of zero values. This posed a unique problem in our training setup: a zeroed K channel in the input data could nullify both the input signal and its gradients when surrogate gradients are used for backpropagation in SNNs. Consequently, this would undermine the learning process by reducing the effectiveness of gradient-based updates, complicating model training and ultimately impacting the accuracy of our results.
Appendix B Tranning progession details
This section contain the training progression graph for each variant of experiments in Section V-B3 and Section V-B1.
It is noteworthy from Figure 4 and 5 that performance fluctuations were observed in both traditional coding and our coding method. We hypothesize that these fluctuations might related to inherent error rate during sampling process in Poisson-distributed spike trains[3].
Another hypothesis is that although bit planes may possess a certain degree of spatial information, they can also encompass noise on low order bit planes that interferes with the training process or have little to none distinguishable features (which might explain how on low noises single-channel image dataset such as MNIST, the performance of only using bit planes coding remain competitive to traditional rate coding with less amount of time step). Consequently, the selective identification of bit planes that retain valuable spatial information continues to pose a significant challenge.
(a) MNIST
(b) Fashion-MNIST
(c) KMNIST
Figure 3: Comparison in accuracy progression and loss progression between coding methods on different low-resolution gray-scale image classification datasets.
(a) CIFAR10[26]
(b) CIFAR100[26]
(c) Caltech101[48]
(d) Caltech256[49]
(e) EuroSAT[50]
(f) Imagenette[51]
(g) Food101[52]
(h) GTSRB[53]
Figure 4: Comparison in loss progression of traditional rate coding with our method on image classification datasets. On every sub-figure, loss details of every color models is shown on the left and average result of all models on the right with maximum loss difference denoted on top of figure.
(a) CIFAR10[26]
(b) CIFAR100[26]
(c) Caltech101[48]
(d) Caltech256[49]
(e) EuroSAT[50]
(f) Imagenette[51]
(g) Food101[52]
(h) GTSRB[53]
Figure 5: Comparison in accuracy progression of traditional rate coding with our method on image classification datasets. On every sub-figure, accuracy details of every color models is shown on the left and average result of all models on the right with maximum accuracy difference denoted on top of figure.